# Plasmon Mode Engineering with Electrons on Helium

### v1

@@
Abstract
@@
An ensemble of electrons trapped above the surface of superfluid helium is a paradigm system for investigating, and controlling, the collective charge dynamics of low-dimensional electronic matter. Of particular interest is the ability to engineer the spatial and spectral structure of surface plasmon modes in this system for integration into hybrid quantum systems or circuit quantum electrodynamic device architectures. Here we present experiments on a hybrid electron-on-helium microchannel device designed to host microwave-frequency plasmon modes having a spatial structure dictated by the geometry of the microchannel confinement. The plasma oscillations are generated via local microwave frequency excitation of the electrons in the microchannel. When this excitation is resonant with a particular surface plasmon-mode it produces a non-equilibrium decrease in the electron conductance, which we detect via simultaneous transport measurements. We find that the spatial structure of the surface plasmons is in excellent agreement with our device design parameters and modeling, and their frequency can be tuned over a broad range (several GHz) by precisely varying the areal density of electrons in the channel. By measuring the plasma resonance spectrum lineshape, and its power dependence, we can quantify the level of spatial homogeneity associated with each plasmon mode. The results highlight the versatility of electrons on helium as a model system for investigating, and engineering, the collective mode structure of low-dimensional Coulomb liquid and solid states and demonstrate a viable path for integrating precisely engineered surface plasmons in electrons on helium with future hybrid circuit quantum electrodynamic systems.
preprint: APS/123-QED
Appendix A Introduction
@@
Circuit quantum electrodynamics (cQED) [1, 2] and the toolkit of microwave frequency optomechanics [3] have enabled the development of sophisticated quantum control and measurement protocols for a wide variety of quantum systems ranging from superconducting circuits [4] and semiconductor spins [5], to systems of trapped electrons [6, 7, 8], as well as nano- and micromechanical oscillators [9, 10, 11, 12]. These techniques can also be leveraged as powerful experimental tools for investigating microwave frequency collective phenomena in quantum systems composed of many interacting particles or degrees of freedom [13]. For example, when coupled with superconducting circuits, these approaches have been used to study collective modes in magnonic [14, 15, 16, 17, 18, 19] and phononic [20, 21, 22, 23, 24, 25] systems and to investigate the dynamics of spin ensembles [26, 27].
Electrons trapped above the surface of condensed noble gas substrates, such as superfluid helium or solid neon, are emerging as promising systems for integration with cQED architectures and microwave frequency devices for quantum information processing [28, 29, 30, 31, 32]. At the level of single electrons, cQED techniques have been used to investigate the in-plane orbital states of electrons on helium [33, 6] and have recently been utilized to realize high-coherence charge qubits on the surface of solidified neon [7, 8]. In contrast to single electron dynamics, these systems can also host a wide variety of collective charge modes including plasmonic [34, 35] and magneto-plasmonic excitations [36, 37, 38, 39, 40], as well as hybrid modes coupling the dynamics of multiple degrees of freedom [41, 42, 43]. Additionally, ensembles of electrons on helium have been strongly coupled to three-dimensional microwave cavities to study cyclotron resonance [44, 45] and integrated into hybrid circuits in which an electron ensemble is placed above a planar microwave resonator [46].
@@
Fully leveraging cQED-type techniques to study the high-frequency dynamics of electrons on helium requires the development of devices that have not only an optimized microwave environment [46], but also the ability to engineer and manipulate the collective modes of the electron system via precise spatial control. In this work, we address the latter of these aspects by realizing a device that enables precision control over the spatial distribution of electrons in a microchannel geometry, providing the ability to engineer, excite, and detect plasmonic excitations with frequencies in a range compatible with cQED-based systems. Local microwave excitation resonantly couples to the plasmon modes, which we detect via changes in the electron conductance determined by simultaneous transport measurements. By precisely varying the electron density in the microchannel, we can tune the frequency of the modes by several GHz. Analyzing the power dependent plasmon mode lineshape allows us to infer the level of spatial homogeneity of electrons in the microchannel, which we compare with simulations of the electron density. Finally, we highlight how this type of device and our results demonstrate the overall system control necessary to integrate with future low-loss microwave cQED architectures.
Figure 1: (a) False color scanning electron micrograph of the microchannel device. Grounded guard electrodes (green) patterned above a resist layer surround reservoir electrodes (yellow) located beneath. The central microchannel region between the two reservoirs consists of a channel electrode (red) and top side gate electrode (blue). An ac drive voltage $V_{\mathrm{ac}}$ at frequency $f_{\mathrm{ac}}$ drives the electrons (white dots) through the channel and the resulting transport signal $V_{\mathrm{s}}$ is detected via a lock-in amplifier. A $\omega/2\pi$ frequency microwave signal on the gate electrode generates longitudinal plasmons (shown schematically in the $\delta n_{s}$ plot). (b) Top panel: Schematic cross section view across ${x=0}$ of the channel with a fixed dc bias voltage (i.e. no microwave drive) on the side gate electrode $V_{\mathrm{gt}}$. Bottom panel: Electrostatic potential profile, $\phi(x=0,y)$, transverse to the channel (solid red line), with chemical potential $\phi_{e}$ (red line), and the distribution of electron density $n_{s}(y)$ along the $y$-direction (teal line). (c) Transport measurements performed at various values of $V_{\mathrm{gt}}$, showing the three characteristic transport regimes. Here, $V_{\mathrm{res}}=0.9$ V, $V_{\mathrm{ac}}=20$ mV, $f_{\mathrm{ac}}=1.408$ MHz. For a complete device and transport description, see Section B.1 and B.2.
Appendix B Experimental Setup
B.1 Microchannel device for plasmon confinement
@@
Microchannel device architectures, like the one we employ, are widely used to study the effect of geometric confinement on the thermodynamic ground state and transport properties of electrons on helium. Typically in these devices, micron-scale deep channels are filled with superfluid helium via capillary forces and electrons are deposited above the superfluid surface. Metallic electrodes around the channels precisely shape the electrostatic environment to control the spatial distribution of surface state electrons and perform transport experiments [47, 48]. These types of channeled devices have been used to reveal dynamical ordering of two-dimensional [49, 50] and quasi-one-dimensional electron chains [51, 52], and perform ultra-efficient clocking of electrons in microchannel-based CCD arrays [53, 54]. Here we leverage a microchannel architecture to engineer the spatial structure of the two-dimensional electron system in order to host and investigate charge density oscillations, i.e. plasmons.
@@
The high degree of spatial control and confinement over the electrons in the reservoirs and microchannel regions are enabled by voltages ($V_{i}$) applied to the four electrodes: $V_{\mathrm{gt}}$ (gate), $V_{\mathrm{ch}}$ (channel), $V_{\mathrm{res}}$ (reservoirs), and the guard electrode. These voltages allow us to control the two-dimensional electrostatic environment experienced by the electrons in the plane of the helium surface $\phi(x,y)=\sum_{i}V_{i}\alpha_{i}(x,y)$ (see Fig. 1b), where the constant $\alpha_{i}(x,y)$ describes the capacitive coupling between the electrons and the corresponding $i^{\mathrm{th}}$ electrode. To design a given confinement profile, we numerically solve the Laplace equation using finite element modelling (FEM) techniques [55]. This allows us to extract $\alpha_{i}$ and construct the potential by applying appropriate values of $V_{i}$. This numerical procedure also allows us to calculate the areal electron density $n_{s}(x,y)$ for a given potential.
By controlling the electrostatic environment in this fashion, we can effectively create a resonant cavity for confined plasmonic modes, i.e. oscillations of the charge density $\delta n_{s}$, along the length of the central channel. At the boundaries of the microchannel, electrons are free to enter and exit into the reservoir regions, enforcing charge density nodes at the ends of the channel, as depicted in Fig. 1a. Because the number of electron rows is $\gtrsim 10$ in the density regime in which we investigate plasmons, the electrons in the central channel can be modeled as a two-dimensional sheet of charge defined by the electrostatic confinement produced by the electrode voltages. The long and narrow geometry ($L\gg w$) of the central channel ensures a large separation in frequency between plasmons along, versus perpendicular, to the channel. This allows us to consider only longitudinal plasmon standing-waves along the channel length, which have the following dispersion relation [34, 56, 57],
@@
where $\omega_{p}$ is the density-dependent frequency of a plasmon having wavevector $q_{x}=n\pi/L$ and mode number $n$, $m_{e}$ is the electron mass, $e$ is the electron charge, and $\varepsilon_{0}$ is the vacuum permittivity. The wavevector-dependent factor $F(q_{x})$ takes into account the reduction in the electron-electron interaction due to the presence of the nearby metallic electrodes. This screening factor will correspondingly reduce the plasmon frequency [34], and for the geometry of our device we utilize the following phenomenological form for $F(q_{x})$,
$F(q_{x})=\frac{1}{2}(\tanh{q_{x}l}+\tanh{q_{x}h}),$ (2)
@@
where $l=w-w_{e}$ parameterizes the effective distance of the electron sheet from the surrounding side gate electrodes and $h=1.4~{}\upmu$m is the height of the electrons above the bottom channel electrode. The effective width $w_{e}$ of the electron system is defined where the parabolic confinement potential of the channel is equal to the chemical potential $\phi_{e}$, i.e. $\phi(y=w_{e}/2)=\phi_{e}$, as shown in the bottom panel of Fig. 1b, which we extract from FEM. For the specific geometry of our device, which includes laterally defined side gate electrodes, an analytical solution for the screening factor is lacking. However, we find that the phenomenological form presented in Eq. 2 captures to good approximation the screening contributions in the long wavelength limit ($q_{x}h,\ q_{x}l\ll 1$), as well as in the limiting case, in which the screening electrodes are moved infinitely far from the electrons in the channel and the unscreened plasmon dispersion is recovered, i.e. $F(q_{x})=1$. Using Eq. 1 and FEM calculations, we find a fundamental ($n=1$) plasmon mode frequency of $\omega_{p}/2\pi\ simeq1.0$ GHz at $n_{s}\simeq 2.3\times 10^{12}$ m-2. This lowest frequency mode corresponds to a half wavelength standing wave of the time-varying change in density $\delta n_{s}$ along the channel, as shown in the top panel of Fig. 1a. In the following section, we discuss how these modes are generated using an additional microwave drive, and detected using transport techniques.
B.2 Transport measurements & microwave excitation
@@
To characterize the electron system in the central microchannel, and its collective dynamics, we utilize a conventional ac transport measurement scheme [51]. In these measurements, an ac voltage $V_{\mathrm{ac}}$ is superimposed on the left reservoir electrode driving electrons from one reservoir to the other via the central channel at a frequency $f_{\mathrm{ac}}$. The resulting electron transport through the channel is detected from the voltage $V_{\textrm{s}}$ induced on the right reservoir electrode, which we measure using standard phase-sensitive lock-in techniques. As described previously, a dc voltage $V_{\mathrm{ch}}$ applied to the channel electrode controls the population of electrons in the central microchannel. In Fig. 1c we show a standard transport map as we tune the electron density and confinement potential in the central channel. This type of measurement reveals three transport regimes depending on the density of electrons in the central microchannel. In regime I, $\phi(y=0)<\phi_{e}$, and electrons cannot enter the channel from the reservoirs. When $\phi(y=0)=\phi_{e}$ electrons enter the microchannel and form a highly conducting state (regime II) in which the electrons interact weakly with the helium surface, resulting in a large transport signal [49]. At sufficiently high density, the electrons in the microchannel form a low-conductivity Wigner solid (regime III) [58]. These measurements also allow us to calculate the electron density in the central microchannel $n_{s}$ from the potential in the center of the channel $\phi^{0}\equiv\phi(x=0,y=0)$ and the chemical potential $\phi_{e}$, as
$n_{s}=\frac{\varepsilon_{\mathrm{He}}\varepsilon_{0}}{eh}(\phi^{0}-\phi_{e})$ (3)
@@
where, $\varepsilon_{\mathrm{He}}=1.057$ is the dielectric constant of liquid helium. Here, the chemical potential is calculated using ${\phi_{e}=V^{\mathrm{th}}_{\mathrm{ch}}\alpha^{0}_{\mathrm{ch}}}$ and the capacitive coupling constant in the center of the microchannel, ${\alpha^{0}_{\mathrm{ch}}\equiv\alpha_{\mathrm{ch}}(x=0,y=0)\simeq 0.7}$, is obtained through FEM calculations of the device.
@@
To generate plasma oscillations of this electron sheet, we apply a high-frequency signal $\omega/2\pi$ to the gate electrodes located on either side of the microchannel as shown in Fig. 1a. The microwave power modulates the otherwise static confinement potential throughout the microchannel, which leads to a periodic modulation of the effective width $w_{e}$ of the electron system (see top panel Fig. 1b). The periodic modulation of the potential creates charge density oscillations of the electrons in the central microchannel due to the strong Coulomb interaction. The frequency and amplitude of these oscillations are controlled by the frequency of the gate modulation and the power $P$ of the microwave signal respectively.
Appendix C Results & Discussion
C.1 Plasmon detection

### v2

@@
Plasmon mode engineering with electrons on helium
Camille A. Mikolas mikolasc@msu.edu Department of Physics & Astronomy, Michigan State University, East Lansing, MI 48824, USA    Niyaz R. Beysengulov Department of Physics & Astronomy, Michigan State University, East Lansing, MI 48824, USA EeroQ Corporation, Chicago, IL 60651, USA    Austin J. Schleusner Department of Physics & Astronomy, Michigan State University, East Lansing, MI 48824, USA    David G. Rees EeroQ Corporation, Chicago, IL 60651, USA    Camryn Undershute Department of Physics & Astronomy, Michigan State University, East Lansing, MI 48824, USA    Johannes Pollanen pollanen@msu.edu Department of Physics & Astronomy, Michigan State University, East Lansing, MI 48824, USA
@@
An ensemble of electrons trapped above superfluid helium offers a paradigm system for investigating and controlling collective charge dynamics in low-dimensional electronic matter. Of particular interest is the ability to spatially control and engineer surface plasmons for integration with hybrid quantum systems and circuit quantum electrodynamic device architectures. Here we present experiments using an electron-on-helium microchannel device that hosts microwave-frequency plasmons, generated via local microwave excitation in an electrostatically defined central channel. By precisely varying the electron density, we demonstrate tunability of plasmon mode frequencies over several GHz. Additionally, we find that the power dependence of these modes can be used to investigate both homogeneous and inhomogeneous sources of spectral broadening. These results demonstrate the versatility of electrons on helium for probing collective excitations in low-dimensional Coulomb liquids and solids, and demonstrate a path for integrating engineered plasmons in electrons on helium with hybrid circuit quantum electrodynamic systems.
Introduction
@@
Circuit quantum electrodynamics (cQED) [1] has enabled the development of sophisticated quantum control and measurement protocols for a wide variety of quantum systems ranging from superconducting circuits [2] and semiconductor spins [3], to systems of trapped electrons [4, 5, 6], as well as nano- and micromechanical oscillators [7, 8, 9]. These techniques can also be leveraged as powerful experimental tools for investigating microwave frequency collective phenomena in quantum systems composed of many interacting particles or degrees of freedom [10]. For example, when coupled with superconducting circuits, these approaches have been used to study collective modes in magnonic [11, 12, 13, 14] and phononic [15, 16, 17, 18] systems and to investigate the dynamics of spin ensembles [19, 20].
@@
Electrons trapped above the surface of condensed noble gas substrates, such as superfluid helium or solid neon, are emerging as promising systems for integration with cQED architectures and microwave frequency devices for quantum information processing [21, 22, 23, 24]. At the level of single electrons, cQED techniques have been used to investigate the in-plane orbital states of electrons on helium [25, 4] and have recently been utilized to realize high-coherence charge qubits on the surface of solidified neon [5, 6]. In contrast to single electron dynamics, these systems can also host a wide variety of collective charge modes including plasmonic [26, 27] and magneto-plasmonic excitations [28, 29, 30, 31, 32], as well as hybrid modes coupling the dynamics of multiple degrees of freedom [33, 34, 35]. Additionally, ensembles of electrons on helium have been strongly coupled to three-dimensional microwave cavities to study cyclotron resonance [36, 37] and integrated into hybrid circuits in which an electron ensemble is placed above a planar microwave resonator [38]. Fully leveraging cQED-type techniques to study the high-frequency dynamics of electrons on helium requires the development of devices that have not only an optimized microwave environment [38], but also the ability to engineer and manipulate the collective modes of the electron system via precise spatial control. In this work, we address the latter of these aspects by realizing a device that enables precision control over the spatial distribution of electrons in a microchannel geometry, providing the ability to engineer, excite, and detect plasmonic excitations with frequencies in a range compatible with cQED-based systems. Local microwave excitation resonantly couples to the plasmon modes, which we detect via changes in the electron conductance determined by simultaneous transport measurements. By precisely varying the electron density in the microchannel, we can tune the frequency of the modes by several GHz. Analyzing the power dependent plasmon response allows us to investigate possible mechanisms leading to plasmon dephasing and energy loss. Finally, we highlight how this type of device and our results demonstrate the overall system control necessary to integrate with future low-loss microwave cQED architectures.
Results and discussion
Microchannel device for plasmon confinement
@@
Microchannel device architectures, like the one we employ, are widely used to study the effect of geometric confinement on the thermodynamic ground state and transport properties of electrons on helium. Typically in these devices, micron-scale deep channels are filled with superfluid helium via capillary forces and electrons are deposited above the superfluid surface. Metallic electrodes around the channels are used to precisely shape the electrostatic environment to control the spatial distribution of surface state electrons and perform transport experiments [39, 40]. These types of channeled devices have been used to reveal dynamical ordering of two-dimensional [41, 42] and quasi-one-dimensional electron chains [43, 44], and perform ultra-efficient clocking of electrons in microchannel-based CCD arrays [45, 46]. Here we leverage a microchannel architecture to engineer the spatial structure of the two-dimensional electron system in order to host and investigate charge density oscillations, i.e. plasmons.
Figure 1: Microchannel device for plasmon-mode engineering. (a) False color scanning electron micrograph of the microchannel device. Grounded guard electrodes (green) patterned above a resist layer surround reservoir electrodes (yellow) located beneath. The central microchannel region between the two reservoirs consists of a channel electrode (red) and top side gate electrode (blue). An ac voltage $V_{\mathrm{ac}}$ at frequency $f_{\mathrm{ac}}$ drives the electrons (white dots) through the channel and the resulting transport signal $V_{\mathrm{s}}$ is detected via a lock-in amplifier. A microwave signal with a frequency $\omega/2\pi$ is applied to the gate electrode to generate longitudinal plasmons in the central channel (the first mode of which is shown schematically by the $\delta n_{s}$ plot). (b) Top panel: Schematic cross section view across ${x=0}$ of the channel with a fixed dc bias voltage (i.e. no microwave drive) on the side gate electrode $V_{\mathrm{gt}}$. Bottom panel: Electrostatic potential profile, $\phi(x=0,y)$, transverse to the channel (solid red line), with chemical potential $\phi_{e}$ (dashed red line), and the distribution of electron density $n_{s}(y)$ along the $y$-direction (solid blue line). (c) Transport measurements performed at various values of $V_{\mathrm{gt}}$, showing the three characteristic transport regimes. Here, $V_{\mathrm{res}}=0.9$ V, $V_{\mathrm{ac}}=20$ mV, $f_{\mathrm{ac}}=1.408$ MHz. See main text for complete description.
@@
The high degree of spatial control and confinement over the electrons in the reservoirs and microchannel regions are enabled by voltages ($V_{i}$) applied to the four electrodes: $V_{\mathrm{gt}}$ (gate), $V_{\mathrm{ch}}$ (channel), $V_{\mathrm{res}}$ (reservoirs), and the guard electrode. These voltages allow us to control the two-dimensional electrostatic environment experienced by the electrons in the plane of the helium surface $\phi(x,y)=\sum_{i}V_{i}\alpha_{i}(x,y)$ (see Fig. 1b), where the constant $\alpha_{i}(x,y)$ describes the capacitive coupling between the electrons and the corresponding $i^{\mathrm{th}}$ electrode. To design a given confinement profile, we numerically solve the Laplace equation using finite element modeling (FEM) techniques [47]. This allows us to extract $\alpha_{i}$ and construct the potential by applying appropriate values of $V_{i}$. This numerical procedure also allows us to calculate the areal electron density $n_{s}(x,y)$ for a given potential.
By controlling the electrostatic environment in this fashion, we can effectively create a resonant cavity for confined plasmonic modes, i.e. oscillations of the charge density $\delta n_{s}$ along the length of the central channel. At the boundaries of the microchannel, electrons are free to enter and exit into the reservoir regions, enforcing charge density nodes at the ends of the channel, as depicted in Fig. 1a. Because the number of electron rows is $\gtrsim 10$ in the density regime in which we investigate plasmons, the electrons in the central channel can be modeled as a two-dimensional sheet of charge defined by the electrostatic confinement produced by the electrode voltages. The long and narrow geometry ($L\gg w$) of the central channel ensures a large separation in frequency between plasmons along versus perpendicular to the channel. This allows us to consider only longitudinal plasmon standing-waves along the channel length, which have the following dispersion relation [26, 48, 49],
@@
where $\omega_{p}$ is the density-dependent frequency of a plasmon having wavevector $q_{x}=n\pi/L$ and mode number $n$, $m_{e}$ is the electron mass, $e$ is the electron charge, and $\varepsilon_{0}$ is the vacuum permittivity (see Supplementary Information, section I). The wavevector-dependent factor $F(q_{x})$ takes into account the reduction in the electron-electron interaction due to the presence of the nearby metallic electrodes. This screening factor will correspondingly reduce the plasmon frequency [26], and for the geometry of our device we utilize the following phenomenological form for $F(q_{x})$,
$F(q_{x})=\frac{1}{2}(\tanh{q_{x}l}+\tanh{q_{x}h}),$ ) ) , (2)
@@
where $l=w-w_{e}$ parameterizes the effective distance of the electron sheet from the surrounding side gate electrodes and $h=1.4~{}\upmu$m is the height of the electrons above the bottom channel electrode. The effective width $w_{e}$ of the electron system is defined where the parabolic confinement potential of the channel is equal to the chemical potential $\phi_{e}$, i.e. $\phi(y=w_{e}/2)=\phi_{e}$, as shown in the bottom panel of Fig. 1b, which we extract from FEM. For our specific device geometry, which includes laterally defined side gate electrodes, an analytical solution for the screening factor is lacking. However, we find that the phenomenological form presented in Equation (2) captures to good approximation the screening contributions in the long wavelength limit ($q_{x}h,\ q_{x}l\ll 1$), as well as in the limiting case, in which the screening electrodes are moved infinitely far from the electrons in the channel and the unscreened plasmon dispersion is recovered, i.e. $F(q_{x})=1$. Using Equation (1) and FEM calculations, we find a fundamental ($n=1$) plasmon mode frequency of $\omega_{p}/2\pi\simeq 1.0$ GHz at $n_{s}\simeq 2.3\times 10^{12}$ m-2. This lowest frequency mode corresponds to a half-wavelength standing wave of the time-varying change in density $\delta n_{s}$ along the channel, as shown in the top panel of Fig. 1a. In the following section, we discuss how these modes are generated using an additional microwave drive and detected using transport techniques.
Transport measurements & microwave excitation
@@
To characterize the electron system in the central microchannel, and its collective dynamics, we utilize a conventional ac transport measurement scheme [43]. In these measurements, an ac voltage $V_{\mathrm{ac}}$ is superimposed on the left reservoir electrode driving electrons from one reservoir to the other via the central channel at a frequency $f_{\mathrm{ac}}$. The resulting electron transport through the channel is detected from the voltage $V_{\textrm{s}}$ induced on the right reservoir electrode, which we measure using standard phase-sensitive lock-in techniques. As described above, a dc voltage $V_{\mathrm{ch}}$ applied to the channel electrode controls the population of electrons in the central microchannel. In Fig. 1c we show a standard transport map as we tune the electron density and confinement potential in the central channel. This type of measurement reveals three transport regimes depending on the density of electrons in the central microchannel. In regime I, $\phi(y=0)<\phi_{e}$, and electrons cannot enter the channel from the reservoirs. When $\phi(y=0)=\phi_{e}$, the channel threshold voltage $V_{\mathrm{ch}}^{\mathrm{th}}$ condition is met and electrons can enter the microchannel for $V_{\mathrm{ch}}\geq V_{\mathrm{ch}}^{\mathrm{th}}$. In this regime (regime II), the electrons form a highly conducting state in which the electrons interact weakly with the helium surface, resulting in a large transport signal [41]. At sufficiently high density, the electrons in the microchannel form a low-conductivity Wigner solid (regime III) [50]. These measurements also allow us to calculate the electron density in the central microchannel $n_{s}$ from the potential in the center of the channel $\phi^{0}\equiv\phi(x=0,y=0)$ and the chemical potential $\phi_{e}$, as
$n_{s}=\frac{\varepsilon_{\mathrm{He}}\varepsilon_{0}}{eh}(\phi^{0}-\phi_{e}),$ (3)
@@
where $\varepsilon_{\mathrm{He}}=1.057$ is the dielectric constant of liquid helium. Here, the chemical potential is calculated using ${\phi_{e}=V^{\mathrm{th}}_{\mathrm{ch}}\alpha^{0}_{\mathrm{ch}}}$ [51] and the capacitive coupling constant in the center of the microchannel, ${\alpha^{0}_{\mathrm{ch}}\equiv\alpha_{\mathrm{ch}}(x=0,y=0)\simeq 0.7}$, is obtained through FEM calculations of the device [44].
@@
To generate plasma oscillations in the electron sheet, we apply a high-frequency signal onto the gate electrodes located on either side of the microchannel as shown in Fig. 1a. The microwave power modulates the otherwise static confinement potential throughout the microchannel, which leads to a periodic modulation of the effective width $w_{e}$ of the electron system (see top panel Fig. 1b) and creates charge density oscillations of the electrons in the central microchannel due to the strong Coulomb interaction. The frequency and amplitude of these oscillations are controlled by the gate modulation frequency $\omega/2\pi$ and microwave signal power $P$, which is measured from the output of the high-frequency source. This microwave signal is attenuated by an additional 28 dB before entering the cryostat.
Figure 2: Transport detected plasmon resonances. Microchannel transport measurements in the presence of a $\omega/2\pi~{}=~{}5.5$ GHz microwave excitation signal on the gate electrode for increasing values of microwave power $P$. Traces are offset vertically for clarity and color coded with the corresponding $P$. Vertical dashed lines indicate the transition into a low-conductivity Wigner crystal electron state, which increases with increasing $P$. At $P\gtrsim-15$ dBm, resonance-like features appear in regime II, indicating plasmon modes generated along the length of the channel as described in the main text. Measurements were performed at $T=18$ mK, with $V_{\mathrm{gt}}=0$ V, $V_{\mathrm{ac}}=6$ mV, $f_{\mathrm{ac}}=3$ MHz, and $V_{\mathrm{res}}~{}=~{}0.4$ V.
Plasmon detection